OpenStudy (anonymous):

Find the exact value of cos (19pi/12)

OpenStudy (anonymous):

I keep getting this question marked wrong I use cos(5pi/6+3pi/4)=cos5pi/6cos3pi/4+sin5pi/6sin3pi/4 this gets me -sqr3/2*-sqr2/2+1/2*sqr2/2 when I simplify this i get sqr6/4+sqr2/2 or (sqr6 + sqr 2)/4 , but it still gets marked wrong. > Could you show me where the mistake in my work is? Thank you c:

OpenStudy (anonymous):

\[\frac{ \sqrt{3}-1 }{ 2\sqrt{2}}\]

OpenStudy (campbell_st):

the mistake is in your expansion cos(A + B) = cos(A)cos(B) - sin(A)sin(B)

OpenStudy (anonymous):

@campbell_st the sign switches?

OpenStudy (campbell_st):

yep... cos(A + B) = cos(A)cos(B) - sin(A)sin(B) cos(A - B)= cos(A)cos(B) + sin(A)sin(B)

OpenStudy (anonymous):

Thank you! No wonder I kept messing cos up!

OpenStudy (campbell_st):

glad to help